OpenStudy (moss65):

prove there does not exist an onto homomorphism f: Z_n -> Z_(n-1)

OpenStudy (anonymous):

Assume \(f\) is a surjective homomorphism. We have the following. \(f(a^n) = f(a)^n = f(0) = 0\) So \(f(a)^n = 0\) and therefore the order of \(f(a)\) divides \(n\). Why is this a contradiction if assuming \(f\) is surjective?

OpenStudy (moss65):

this something about isomorphism: ord(a) =ord(f(a)), but it cannot be because ord(a)=n and ord(f(a))=n-1

OpenStudy (moss65):

so f cannot be isomorphic, and it is not one-to-one

OpenStudy (anonymous):

Does the order of each element in \(Z_{n -1}\) divide \(n\)?

OpenStudy (anonymous):

We are not talking about an isomorphism here. We only know that \(f\) is surjective.

OpenStudy (moss65):

no, it doesn't divide n

OpenStudy (anonymous):

Well, the order of the identity element (order 1) does divide \(n\). However, since this is a surjection, for each \(b \in Z_{n - 1}\) we can find some \(a \in Z_n\) such that f(a) = b. Therefore it must hold that for each \(b \in Z_{n - 1}\) that \(ord(b) \mid n\).

OpenStudy (anonymous):

Hint: What is the order of the generator in \(Z_{n - 1}\)?

OpenStudy (moss65):

yes, but it cannot be. is this contradiction you meant?

OpenStudy (anonymous):

What is the order of \(1\) in \(Z_{n−1}\)?

OpenStudy (anonymous):

We will get our contradiction if you can answer this question. Does \(|1_{Z_{n-1}}|\) divide \(n\)?

OpenStudy (moss65):

Sorry Alchemista, there was some problem with my comp. The order 1 bar in Z_n-1 is n-1, so it definitely doesn't divide n

OpenStudy (anonymous):

Exactly, and since we demonstrated above that if \(f\) is a surjective homomorphism \(\forall x\in Z_{n−1}[|x| \text{ divides } n]\) we have arrived at our contradiction. Since \(|1|\) does not divide \(n\).

OpenStudy (anonymous):

Does that make sense?

OpenStudy (moss65):

absolutely. Thank you so much. It was nice to work with you again.